jane is 5 years older than tom.
three times jane's age plus twice toms age is 35.
what are the ages of jane and tom.

Answer by ReadingBoosters(3246)

x = jane
y = tom
x = y + 5
3x + 2y = 35
Substitute into the second equation
3(y+5) +2y = 35
3y + 15 + 2y = 35
5y = 35-15
5y = 20
y = 4
x = 4+5=9
Jane is 9 years old
Tom is 4 years old
Proof
4+5 = 9
3(9) + 2(4) = 35